# DietGlance: Dietary Monitoring and Personalized Analysis at a Glance with Knowledge-Empowered AI Assistant

### v2

@@
Abstract
Growing awareness of wellness has prompted people to consider whether their dietary patterns align with their health and fitness goals. In response, researchers have introduced various wearable dietary monitoring systems and dietary assessment approaches. However, these solutions are either limited to identifying foods with simple ingredients or insufficient in providing analysis of individual dietary behaviors with domain-specific knowledge. In this paper, we present DietGlance, a system that automatically monitors dietary in daily routines and delivers personalized analysis from knowledge sources. DietGlance first detects ingestive episodes from multimodal inputs using eyeglasses, capturing privacy-preserving meal images of various dishes being consumed. Based on the inferred food items and consumed quantities from these images, DietGlance further provides nutritional analysis and personalized dietary suggestions, empowered by the retrieval augmentation generation module on a reliable nutrition library. A short-term user study (N=33) and a four-week longitudinal study (N=16) demonstrate the usability and effectiveness of DietGlance, offering insights and implications for future AI-assisted dietary monitoring and personalized healthcare intervention systems using eyewear.
1 Introduction
Dietary behaviors and human health are inextricably linked [73, 47]. For instance, excessive consumption of processed foods high in sugars, unhealthy fats, and sodium increases the risk of chronic diseases [51]. Even small dietary improvements can significantly benefit health [27]. Thus, many individuals have sought dietary monitoring and analysis methods to refine their dietary choices. By offering insights into nutritional intake, these practices enable individuals to make informed decisions on diet planning, correct nutritional imbalances, and achieve health goals such as weight management or disease prevention [2]. The dominant dietary monitoring and analysis method nowadays is food journaling (on paper, websites, or mobile apps [70]), such as MyFitnessPal [54] and Boohee [11]. These tools allow users to log their food consumption, track nutrients, reflect on their dietary habits, and seek dietary recommendations from nutrition experts. However, these methods face several limitations.
@@ 1 Introduction
Fig. 1: The flow of DietGlance. (a) DietGlance uses Aria Glasses to detect ingestive episodes and collects data in real-world, unconstrained environments. Users can view their dietary analysis results using the (b) mobile (from left to right: meal log image presentation, detailed nutritional analysis, and personalized suggestions) or (c) desktop interface.
@@
First, manual food journaling is labor-intensive and time-consuming [17, 58]. Most existing tools, like mobile apps, require users to input information or capture images manually, hindering long-term tracking sustainability [67, 18, 49]. Second, comprehensive nutritional analysis remains insufficiently addressed in current systems, which is essential for evaluating dietary intake and making informed decisions. Mobile apps offer basic nutritional insights with pre-built databases but often lack data on customized or home-cooked meals, leading to vague and inaccurate analysis. They also lack real-time, personalized suggestions tailored to individual contexts. Therefore, a holistic automatic dietary monitoring and analysis system supporting food journaling, nutritional analysis, and dietary suggestions is in great need.
@@
To automate food journaling, researchers have explored eyewear-based devices for detecting ingestive episodes [5, 70, 21]. Eyeglasses are both socially acceptable and conveniently positioned near the eating region, making them ideal for this purpose [70, 5]. Contemporary commercial eyewear devices (e.g., Vision Pro [3] and Meta smart glasses [79]), which integrate advanced sensing capabilities for diverse applications, highlight the potential of eyewear for dietary data collection in daily contexts. However, existing systems are only capable of handling simple dishes and idealized dining environments, which restricts their applicability in real-world settings. Approaches for more generic and sophisticated nutritional analysis remain largely underexplored.
@@
Recent advances in Large Language Models (LLMs) offer new opportunities. Their extensive knowledge and contextual understanding enable effective management of the complexity and diversity of food items, dietary habits, and nutritional needs across various populations [74, 42]. For example, Szymanski et al. [74] collaborated with registered dietitians to validate the GPT-4-generated food product explanation based on product names, nutrition labels, and dietary goals. Lo et al. [42] applied GPT-4V to dietary assessment, highlighting its effectiveness in food detection under challenging conditions. However, directly using LLM may generate plausible-sounding but incorrect or inaccurate results (i.e., hallucination). Moreover, it may show limited and generic health guidelines that are not personalized to individuals’ needs, overlook comprehensive dietary patterns, and miss relevant nutrition information [74, 26, 55]. There is a great research space in enhancing LLMs by domain knowledge for comprehensive nutritional analysis, enabling more reliable and personalized recommendations. Such a system can facilitate the early detection of nutritional issues and support informed dietary decisions to improve personal health.
@@
To address these gaps, we first conducted a needs-finding study with 45 participants to deepen our understanding of user needs and validate limitations identified in prior research. Consequently, we formulated four key design requirements for a holistic dietary monitoring and analysis system: supporting automated accurate diet identification and logging, facilitating nutritional analysis based on reliable sources, providing personalized dietary suggestions with domain-specific knowledge, while protecting privacy. Following these design requirements, we introduced DietGlance (Fig.2), an AI-assisted system for dietary monitoring and analysis using eyewear, which automatically identifies various food and drink items, and further provides knowledge-supported nutritional analysis and personalized suggestions. Developed using Aria Glasses [72], we proposed a multimodal sensing framework for ingestive episode detection and key-moment meal image capture in free-living conditions. Based on the consumed type and amount of food and drink identified from processed meal images using vision foundation models, a Retrieval-Augmented Generation (RAG) module on a nutrition library is built to empower LLM in providing nutrition analysis and personalized dietary suggestions with knowledge sources incorporating individual profiles and meal logs. The system’s user interface allows users to review personalized dietary suggestions with nutritional analysis, interactively correct errors, add information, and access additional insights.
@@
  • A four-week longitudinal study (N=16) highlighting the positive long-term effects of DietGlance on dietary behaviors. The quantitative analysis of nutrition consumption trends and qualitative analysis of participants’ feedback demonstrates that using DietGlance increased the understanding of their diet, raised their awareness of regular meal timing, promoted healthier dietary habits, and reduced consumption of unhealthy foods.
@@
2 Related Work
2.1 Ingestive Episode Detection
@@
Ingestive episode detection is essential for automatic food journaling [76]. Various wearable sensors have been explored for automated tracking. In [28], Guan et al. developed a multi-source domain adaptation method to capture chewing activity for food type recognition utilizing IMU from wearable. Wrist-mounted devices can track hand movements during eating [19, 33, 78], but these can be distracted by other gestures and require use on the dominant hand. More reliable approaches involve head- or neck-mounted devices like necklace [87], headbands [7], caps [89], and ear-mounted sensors [6, 88, 8], though they are often uncomfortable and socially intrusive. In contrast, eyeglasses offer a more acceptable alternative, with their proximity to the mouth allowing for better detection of eating episodes. Previous work using eyeglasses includes sensors such as electromyography [86], piezoelectric [21], load cells [16], and acoustic sensors [46] for detection of ingestive behavior. Rahman et al. [61] utilized the Inertial Measurement Unit (IMU) in Google Glass for detection. IMU sensors combine accelerometers, gyroscopes, and occasionally magnetometers to measure the specific force, angular rate, and sometimes magnetic field of objects, enabling the tracking of motion and orientation in three-dimensional space [5]. Systems like FitByte [5] integrated IMU and proximity sensors, while MyDJ [70] utilized piezoelectric and accelerometer sensors.
@@
However, prior studies were limited to simple dishes or controlled settings and often relied on crowd-sourced efforts. DietGlance introduces a multimodal sensing framework to record the meal throughout dining in uncontrolled environments of the real world, allowing comprehensive nutritional analysis and personalized diet suggestions.
2.2 Diet Identification and Nutritional Analysis
@@
Alternative methods include uploading meal photos for expert evaluation [48] or crowdsourced nutrient estimation [56], which are labor-intensive and time-consuming, limiting the system’s accessibility and scalability. Additionally, online searches or forums often provide irrelevant information, reducing their effectiveness. With advances in computer vision, some solutions perform food segmentation, volume estimation, and database-based nutrient calculations sequentially [50, 59, 45, 43], while end-to-end neural networks directly estimate nutritional content from meal images [29, 20, 44]. Although they have achieved promising results with pre-built nutrition estimation datasets, the generalization of models limits their applicability to open-world settings, the close setting to daily routines, and vision models alone lack comprehensive knowledge of diet or nutrition [80].
@@
The rise of LLMs offers new potential. They excel in knowledge integration and cross-domain generalization, attributed to their immense parameters and diverse, extensive training data [10]. Lo et al. [42] demonstrated that GPT-4V can identify food items with high accuracy, deduce portion sizes of eating consumption at a comparable performance to dietitians’ estimates, and estimate nutritional components aligning with the USDA National Nutrient Database [^1] USDA National Nutrient Database: https://fdc.nal.usda.gov/index.html under semi free-living conditions. Building on this, our DietGlance system employs GPT-4V with crafted prompt and user profile to estimate diet type and amount. Unlike prior work, we proposed a multimodal sensing framework combining IMU, audio, and images for automatic diet identification and nutritional analysis under completely free-living conditions. Insights from our user studies highlight opportunities to further refine dietary monitoring and analysis systems.
2.3 Personalized Dietary Suggestions
@@
Recent studies have explored machine learning models for personalized health and disease management suggestions. For example, Mitchell et al. [24] developed GlucoGoalie, which combines machine learning and a rule-based expert system to generate dietary goals for individuals with type 2 diabetes. To enhance flexibility and generalization, researchers have turned to LLMs for dietary guidance, leveraging their extensive knowledge and reasoning capabilities. Chatelan et al. [14] investigated the ability of ChatGPT to provide nutritional and dietary guidance for patients with type 2 diabetes and hemodialysis, and their findings show that the generated diet recipes are in accordance with the Diabetes Plate Method. Ataguba et al. [4] explored using LLM for personalized recipe generation and weight-loss management. Further, Yang et al. [84] introduced ChatDiet, an LLM-powered framework for personalized nutrition-oriented food recommender chatbots, and evaluated it on a case study including one-person information, where the generated food recommendation dialogue demonstrates the explainability and personalization. However, researchers noted that LLMs are easily struggling with hallucinations. One of study in [74] conducted an analysis involving registered dietitians to assess the capabilities of LLMs in providing nutritionally accurate and personalized information, showing the outputs of some instances do not align with standards upheld by registered dietitians and contain falsehoods, which can be misleading. Similarly, in [26], researchers found that while the nutritional advice proposed by ChatGPT is generally accurate, it has the potential to produce unsafe diets containing allergens.
@@
2.4 Leveraging RAG to Enhance LLMs for Accurate Knowledge Retrieval
LLMs are widely used for information extraction and summarization due to their vast knowledge and reasoning capabilities. However, directly prompting LLMs can lead to plausible-sounding but inaccurate answers, known as hallucinations, and a lack of domain-specific knowledge. To address these issues, RAG enhances LLMs by integrating a retrievable memory that incorporates knowledge from external sources [30, 25]. For example, Fok et al. [23] use RAG to expand abstracts with additional information from full papers, while Zulfkar et al. [90] develop Memoro to infer the user’s memory needs in conversation and presenting suggestions queried from memories using a RAG module. The Arizona Water Chatbot [62] employs RAG to retrieve water-related information from reputable sources, improving decision-making. Ren et al. [85] explore memory retrieval and generation refinement using RAG for enhanced conversational AI, and Yang et al. [83] develop AQuA that combines software UI elements associated with questions as the query and generates answers using RAG-powered GPT-4 from official documentation and tutorial resources. While RAG has proven effective in handling large knowledge repositories, its application in integrating dietary data has not yet been comprehensively explored.
3 Existing Practices of Dietary Monitoring and Analysis

### v3

@@
Abstract.
Growing awareness of wellness has prompted people to consider whether their dietary patterns align with their health and fitness goals. In response, researchers have introduced various wearable dietary monitoring systems and dietary assessment approaches. However, these solutions are either limited to identifying foods with simple ingredients or insufficient in providing an analysis of individual dietary behaviors with domain-specific knowledge. In this paper, we present DietGlance, a system that automatically monitors dietary behaviors in daily routines and delivers personalized analysis from knowledge sources. DietGlance first detects ingestive episodes from multimodal inputs using eyeglasses, capturing privacy-preserving meal images of various dishes being consumed. Based on the inferred food items and consumed quantities from these images, DietGlance further provides nutritional analysis and personalized dietary suggestions, empowered by the retrieval-augmented generation module on a reliable nutrition library. A short-term user study (N=33) and a four-week longitudinal study (N=16) demonstrate the usability and effectiveness of DietGlance, offering insights and implications for future AI-assisted dietary monitoring and personalized healthcare intervention systems using eyewear.
Eating Detection, Dietary Monitoring, Nutritional Analysis, Multimodal Sensing
copyright: acmcopyrightjournalyear: 0doi: 0journal: HEALTHjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0ccs: Human-centered computing Ubiquitous and mobile computingccs: Applied computing Health informaticsccs: Computing methodologies Artificial intelligence
1. Introduction
Dietary behaviors and human health are inextricably linked (Stylianou et al., 2021; Marshall et al., 2022). For instance, excessive consumption of processed foods high in sugars, unhealthy fats, and sodium increases the risk of chronic diseases (Micha et al., 2017). Even small dietary improvements can significantly benefit health (Grummon et al., 2023). Thus, many individuals have sought dietary monitoring and analysis methods to refine their dietary choices. By offering insights into nutritional intake, these practices enable individuals to make informed decisions on diet planning, correct nutritional imbalances, and achieve health goals such as weight management or disease prevention (Adams et al., 2020). The dominant dietary monitoring and analysis method nowadays is food journaling (on paper, websites, or mobile apps (Shin et al., 2022)), such as MyFitnessPal (MyFitnessPal, 2024) and Boohee (Boohee, 2024). These tools allow users to log their food consumption, track nutrients, reflect on their dietary habits, and seek dietary recommendations from nutrition experts. However, these methods face several limitations.
@@ 1. Introduction
Fig. 1. The flow of DietGlance. (a) DietGlance uses Aria Glasses to detect ingestive episodes and collects data in real-world, unconstrained environments. Users can view their dietary analysis results using the (b) mobile (from left to right: meal log image presentation, detailed nutritional analysis, and personalized suggestions) or (c) desktop interface.
@@
First, manual food journaling is labor-intensive and time-consuming (Cordeiro et al., 2015a; Päßler et al., 2011). Most existing tools, like mobile apps, require users to input information or capture images manually, hindering long-term tracking sustainability (Schoeller, 1995; Cordeiro et al., 2015b; Merck et al., 2016). Second, comprehensive nutritional analysis remains insufficiently addressed in current systems, which is essential for evaluating dietary intake and making informed decisions. Mobile apps offer basic nutritional insights with pre-built databases but often lack data on customized or home-cooked meals, leading to vague and inaccurate analysis. They also lack real-time, personalized suggestions tailored to individual contexts. Therefore, a holistic automatic dietary monitoring and analysis system supporting food journaling, nutritional analysis, and dietary suggestions is in great need.
@@
To automate food journaling, researchers have explored eyewear-based devices for detecting ingestive episodes (Bedri et al., 2020; Shin et al., 2022; Farooq and Sazonov, 2017). Eyeglasses are both socially acceptable and conveniently positioned near the eating region, making them ideal for this purpose (Shin et al., 2022; Bedri et al., 2020). Contemporary commercial eyewear devices (e.g., Vision Pro (Apple Inc., 2024) and Meta smart glasses (Waisberg et al., 2024)), which integrate advanced sensing capabilities for diverse applications, highlight the potential of eyewear for dietary data collection in daily contexts. However, existing systems are only capable of handling simple dishes and idealized dining environments, which restricts their applicability in real-world settings. Approaches for more generic and sophisticated nutritional analysis remain largely underexplored.
@@
Recent advances in Large Language Models (LLMs) offer new opportunities. Their extensive knowledge and contextual understanding enable effective management of the complexity and diversity of food items, dietary habits, and nutritional needs across various populations (Szymanski et al., 2024; Lo et al., 2024). For example, Szymanski et al. (Szymanski et al., 2024) collaborated with registered dietitians to validate the GPT-4-generated food product explanation based on product names, nutrition labels, and dietary goals. Lo et al. (Lo et al., 2024) applied GPT-4V to dietary assessment, highlighting its effectiveness in food detection under challenging conditions. However, directly using LLM may generate plausible-sounding but incorrect or inaccurate results (i.e., hallucination). Moreover, it may show limited and generic health guidelines that are not personalized to individuals’ needs, overlook comprehensive dietary patterns, and miss relevant nutrition information (Szymanski et al., 2024; Garcia, 2023; Niszczota and Rybicka, 2023). There is significant potential in enhancing LLMs by domain knowledge for comprehensive nutritional analysis, enabling more reliable and personalized recommendations. Such a system can facilitate the early detection of nutritional issues and support informed dietary decisions to improve personal health.
Fig. 2. System Overview.
@@
To address these gaps, we first conducted a needs-finding study with 45 participants to deepen our understanding of user needs and validate limitations identified in prior research. Consequently, we formulated four key design requirements for a holistic dietary monitoring and analysis system: supporting automated, accurate diet identification and logging, facilitating nutritional analysis based on reliable sources, providing personalized dietary suggestions with domain-specific knowledge, while protecting privacy. Following these design requirements, we introduced DietGlance (Fig.2), an AI-assisted system for dietary monitoring and analysis using eyewear, which automatically identifies various food and drink items, and further provides knowledge-supported nutritional analysis and personalized suggestions. Developed using Aria Glasses (Somasundaram et al., 2023), we proposed a multimodal sensing framework for ingestive episode detection and key-moment meal image capture in free-living conditions. Based on the consumed type and amount of food and drink identified from processed meal images using vision foundation models, a Retrieval-Augmented Generation (RAG) module on a nutrition library is built to empower LLM in providing nutrition analysis and personalized dietary suggestions with knowledge sources incorporating individual profiles and meal logs. The system’s user interface allows users to review personalized dietary suggestions with nutritional analysis, interactively correct errors, add information, and access additional insights.
@@
  3. (3)
    A four-week longitudinal study (N=16) highlighting the positive long-term effects of DietGlance on dietary behaviors. The quantitative analysis of nutrition consumption trends and qualitative analysis of participants’ feedback demonstrate that using DietGlance increased the understanding of their diet, raised their awareness of regular meal timing, promoted healthier dietary habits, and reduced consumption of unhealthy foods.
  4. (4)
@@
2. Related Work
2.1. Ingestive Episode Detection
@@
Ingestive episode detection is essential for automatic food journaling (Tang and Hoover, 2025). Various wearable sensors have been explored for automated tracking. Wrist-mounted devices can track hand movements during eating (Dong et al., 2014; Kyritsis et al., 2019; Thomaz et al., 2015), but these can be distracted by other gestures and require use on the dominant hand. More reliable approaches involve head- or neck-mounted devices like necklace (Zhang et al., 2020), headbands (Bi et al., 2017), caps (Zhou and Lukowicz, 2020), and ear-mounted sensors (Bedri et al., 2017; Zhang et al., 2024; Bi et al., 2018), though they are often uncomfortable and socially intrusive. In contrast, eyeglasses offer a more acceptable alternative, with their proximity to the mouth allowing for better detection of eating episodes. Previous work using eyeglasses includes sensors such as electromyography (Zhang and Amft, 2018), piezoelectric (Farooq and Sazonov, 2017), load cells (Chung et al., 2017), and acoustic sensors (Mahmud et al., 2024) for the detection of ingestive behavior. Among these sensors, the Inertial Measurement Unit (IMU) is a widely adopted choice, primarily because it is low-cost, ubiquitous in consumer electronics, and highly effective at capturing the specific 3D motion and orientation of the head during ingestive behaviors. Rahman et al. (Rahman et al., 2015) utilized the IMU in Google Glass for detection. IMU sensors combine accelerometers, gyroscopes, and occasionally magnetometers to measure the specific force, angular rate, and sometimes magnetic field of objects, enabling the tracking of motion and orientation in three-dimensional space (Bedri et al., 2020). Systems like FitByte (Bedri et al., 2020) integrated IMU and proximity sensors, while MyDJ (Shin et al., 2022) utilized piezoelectric and accelerometer sensors. Similarly, other work has explored using deep learning with optical tracking sensors in smart glasses to monitor eating behavior (Stankoski et al., 2024).
@@
However, prior studies were limited to simple dishes or controlled settings (Shin et al., 2022; Chen et al., 2023) and often relied on crowd-sourced efforts (Bedri et al., 2020). DietGlance introduces a multimodal sensing framework to record the meal throughout dining in uncontrolled environments of the real world, allowing comprehensive nutritional analysis and personalized diet suggestions.
2.2. Diet Identification and Nutritional Analysis
@@
While the prior works in Sec. 2.1 focused primarily on detecting ingestive episodes, their methods for diet identification were often simple and used in controlled settings (Shin et al., 2022; Chen et al., 2023; Bedri et al., 2020). This section, therefore, reviews more dedicated approaches for diet identification and nutritional analysis. Traditional expert- or crowd-based methods include uploading meal photos for expert evaluation (Martin et al., 2008) or crowdsourced nutrient estimation (Noronha et al., 2011), which are labor-intensive and time-consuming, limiting the system’s accessibility and scalability. Additionally, online searches or forums often provide irrelevant information, reducing their effectiveness. With advances in computer vision, some solutions perform food segmentation, volume estimation, and database-based nutrient calculations sequentially (Meyers et al., 2015; Pouladzadeh et al., 2014; Lu et al., 2021; Lo et al., 2020), while end-to-end neural networks directly estimate nutritional content from meal images (He et al., 2021; Ege and Yanai, 2017; Lu et al., 2018). Although they have achieved promising results with pre-built nutrition estimation datasets, the generalization of models limits their applicability to open-world settings, and vision models alone lack comprehensive knowledge of diet or nutrition (Wang et al., 2022).
@@
The rise of LLMs offers new potential. They excel in knowledge integration and cross-domain generalization, attributed to their immense parameters and diverse, extensive training data (Bommasani et al., 2021). Lo et al. (Lo et al., 2024) demonstrated that GPT-4V can identify food items with high accuracy, deduce portion sizes of eating consumption at a comparable performance to dietitians’ estimates, and estimate nutritional components aligning with the USDA National Nutrient Database (Haytowitz et al., 2018) under a controlled study protocol that evaluated GPT-4V on pre-collected meal images. Building on this, our DietGlance system employs GPT-4V with crafted prompts and user profiles to estimate diet type and amount. Unlike prior work, we proposed a multimodal sensing framework combining IMU, audio, and images for automatic diet identification and nutritional analysis under completely free-living conditions. Insights from our user studies highlight opportunities to further refine dietary monitoring and analysis systems.
2.3. Personalized Dietary Suggestions
@@
The use of machine learning models for personalized nutrition and dietary recommendation has been widely explored, as documented in several systematic reviews (Varshney et al., 2023; Tsolakidis et al., 2024), marking a shift away from the traditional “one-diet-fits-all” era (Roman et al., 2024). For example, Mitchell et al. (G. Mitchell et al., 2021) developed GlucoGoalie, which combines machine learning and a rule-based expert system to generate dietary goals for individuals with type 2 diabetes. To enhance flexibility and generalization, researchers have turned to LLMs for dietary guidance, leveraging their extensive knowledge and reasoning capabilities. Chatelan et al. (Chatelan et al., 2023) investigated the ability of ChatGPT to provide nutritional and dietary guidance for patients with type 2 diabetes and hemodialysis, and their findings show that the generated diet recipes are in accordance with the Diabetes Plate Method. Ataguba et al. (Ataguba and Orji, 2025) explored using LLM for personalized recipe generation and weight-loss management. Further, Yang et al. (Yang et al., 2024b) introduced ChatDiet, an LLM-powered framework for personalized nutrition-oriented food recommender chatbots, and evaluated it on a case study including personal information, where the generated food recommendation dialogue demonstrates the explainability and personalization. However, researchers noted that LLMs remain susceptible to hallucinations. A study by Szymanski et al. (Szymanski et al., 2024) involved registered dietitians in assessing the capabilities of LLMs in providing nutritionally accurate and personalized information. The results showed that the outputs often failed to align with professional standards and contained falsehoods, which could be misleading. Similarly, in (Garcia, 2023), researchers found that while the nutritional advice proposed by ChatGPT is generally accurate, it has the potential to produce unsafe diets containing allergens.
@@
2.4. Leveraging RAG to Enhance LLMs for Accurate Knowledge Retrieval
LLMs are widely used for information extraction and summarization due to their vast knowledge and reasoning capabilities. However, directly prompting LLMs can lead to plausible-sounding but inaccurate answers, known as hallucinations, and a lack of domain-specific knowledge. To address these issues, RAG enhances LLMs by integrating a retrievable memory that incorporates knowledge from external sources (Huang et al., 2024; Gao et al., 2023). For example, Fok et al. (Fok et al., 2024) use RAG to expand abstracts with additional information from full papers, while Zulfkar et al. (Zulfikar et al., 2024) develop Memoro to infer the user’s memory needs in conversation and present suggestions queried from memories using a RAG module. The Arizona Water Chatbot (Rajan et al., 2024) employs RAG to retrieve water-related information from reputable sources, improving decision-making. Ren et al. (Yen and Zhao, 2024) explore memory retrieval and generation refinement using RAG for enhanced conversational AI, and Yang et al. (Yang et al., 2024a) develop AQuA that combines software UI elements associated with questions as the query and generates answers using RAG-powered GPT-4 from official documentation and tutorial resources. Recent work has specifically explored RAG in specialized health domains to address the accuracy limitations of baseline LLMs. For instance, in (Bingöl et al., 2025), RAG was used to improve the accuracy of dietary principle determination in patients with Chronic Kidney Disease (CKD), showing that RAG generated more accurate and stable results compared to LLMs alone. Similarly, Bano et al. (Bano et al., 2024) found that the augmented model generated more concise and easily comprehensible responses, which was critical for improving patient health literacy. This recent work has shown the promise of RAG in specialized, offline nutritional applications. In this work, we integrate the RAG module within an end-to-end system that automatically detects and analyzes meals in completely free-living conditions, and investigate the performance with and without RAG.
3. Existing Practices of Dietary Monitoring and Analysis